OpenStudy (anonymous):

a 2.0x 10^3 kg car accelerates from rest under the actions of two forces. One is a forward force of1140 N provided by traction between the wheel and the road. The other is 950 N resistive force due to various frictional forces. Use the work-kinetic energy theorem to determine how for the car must travel for its speed to reach 2.0m/s.

OpenStudy (anonymous):

On travelling x meters W=190*x=change in KE=4*10^3